OpenStudy (anonymous):

Please Help With 3 Problems!

OpenStudy (anonymous):

Do you mean x?

OpenStudy (anonymous):

No y

OpenStudy (anonymous):



OpenStudy (anonymous):

There is no y...

OpenStudy (anonymous):

No you solve for Y

OpenStudy (anonymous):

Thats how the problem is written

OpenStudy (anonymous):

What is the equation of the line expressed in slope-intercept form? A. y = –3x – 1 B. y = 3x + 2 C. y = 2x + 3 D. y = –2x + 3

OpenStudy (anonymous):

They probably meant x, and for the second question you want it in y=mx+b form where m is the slope and b is the y - intercept in other words where the line crosses the y - axis.

OpenStudy (anonymous):

So what is the y intercept

OpenStudy (anonymous):

For 2?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Doesnt give me one...

OpenStudy (anonymous):

These are how the problems are written lol (horribly)

OpenStudy (anonymous):

Number 2 I couldnt find the Y-Intercept form thats why I asked for help...

OpenStudy (anonymous):